# Lee–Yang Zeros and Pseudocritical Drift in J-Q Néel–VBS Transitions

### 16 Mar 2026 v1

@@ Abstract
The CBJQ transition provides a controlled first-order benchmark for our Lee–Yang diagnostics [74, 59, 35]. As shown in Fig. 3, we examine two nearby couplings around the literature estimate $J_{c}/Q=0.2175\pm 0.0001$ [74]. At $J/Q=0.2174$, the leading zeros in both the Néel and PSS channels drift clearly toward the first-order spacetime-volume law $g_{\mathrm{LY}}(L)\sim L^{-(d+z)}$ in the ground-state window $\beta\propto L^{z}$ [44, 2, 7, 3, 4, 5], corresponding to $y_{h}=d+z$ and hence $\eta\to-1$ for $d=2$, $z=1$ [24, 26, 30]. The paired-size flow is likewise well described by the first-order reference form $\eta_{\mathrm{eff}}(L)=\eta+c_{1}L^{-\omega}+c_{2}L^{-2\omega}$ with fixed $\eta=-1$ [44, 2, 7, 3, 4]. At $J/Q=0.2175$, the Néel-channel zeros remain broadly consistent with this trend, whereas in the PSS channel the first and second zeros split visibly in their finite-size flow. We interpret this as indicating that $J/Q=0.2175$ lies slightly on the Néel side of the transition, so that a small PSS-channel Lee–Yang edge remains finite on the accessible sizes: the first zero reaches the edge-controlled regime earlier than the second, and the corresponding $\eta_{\mathrm{eff}}(L)$ values should therefore not be viewed as following a single critical scaling form [19, 24, 1, 45]. By contrast, the $J/Q=0.2174$ data provide the cleaner first-order benchmark in both channels. Overall, the CBJQ results establish the expected first-order Lee–Yang phenomenology and highlight the sensitivity of Lee–Yang zeros to coexistence versus critical scaling, even within a very narrow coupling window [74, 59, 44, 3, 4]. This provides a useful first-order baseline for the J-Q cases discussed below [53, 58, 74, 39, 62].
@@
Acknowledgments. We thank Q. Ye, X. Li, J. Rong and W. Zhu for helpful discussions. Z.W. acknowledges support from the China Postdoctoral Science Foundation under Grant No. 2024M752898. This project is supported by the Scientific Research Project (No. WU2025B011) and the Start-up Funding of Westlake University. H.Z. is supported by the National Natural Science Foundation of China (Grant No. 12274126). The authors thank the IT service office and the high-performance computing center of Westlake University.

### 30 Mar 2026 v2

@@ Abstract
The CBJQ transition provides a controlled first-order benchmark for our Lee–Yang diagnostics [105, 80, 44]. As shown in Fig. 3, we examine two nearby couplings around the literature estimate $J_{c}/Q=0.2175\pm 0.0001$ [105]. At $J/Q=0.2174$, the leading zeros in both the Néel and PSS channels drift clearly toward the first-order spacetime-volume law $g_{\mathrm{LY}}(L)\sim L^{-(d+z)}$ in the ground-state window $\beta\propto L^{z}$ [62, 5, 9, 6, 7, 8], implying $y_{h}=d+z$ and hence $\eta\to-1$ for $d=2$, $z=1$ [32, 34, 38]. The paired-size flow is likewise well described by the first-order reference form $\eta_{\mathrm{eff}}(L)=\eta+c_{1}L^{-\omega}+c_{2}L^{-2\omega}$ with fixed $\eta=-1$ [62, 5, 9, 6, 7]. At $J/Q=0.2175$, the Néel-channel zeros remain broadly consistent with this trend, while in the PSS channel the first and second zeros split visibly. We interpret this as indicating that $J/Q=0.2175$ lies slightly on the Néel side of the transition, so that a small PSS-channel Lee–Yang edge remains finite on the accessible sizes: the first zero enters the edge-controlled regime earlier than the second, and the corresponding $\eta_{\mathrm{eff}}(L)$ values should therefore not be viewed as following a single critical scaling form [25, 32, 4, 64]. The $J/Q=0.2174$ data thus provide the cleaner first-order benchmark in both channels. Overall, the CBJQ results establish the expected first-order Lee–Yang phenomenology and highlight the sensitivity of Lee–Yang zeros to coexistence versus critical scaling even within a narrow coupling window [105, 80, 62, 6, 7]. This provides a useful first-order baseline for the J-Q cases discussed below [72, 78, 105, 56, 84].
@@
Acknowledgments. We thank Q. Ye, X. Li, J. Rong and W. Zhu for helpful discussions. Z.W. acknowledges support from the China Postdoctoral Science Foundation under Grant No. 2024M752898. Z.L. is supported by the China Postdoctoral Science Foundation under Grants No. 2024M762935. This project is supported by the Scientific Research Project (No. WU2025B011) and the Start-up Funding of Westlake University. H.Z. is supported by the National Natural Science Foundation of China (Grant No. 12274126). The authors thank the IT service office and the high-performance computing center of Westlake University.